# VLM See, Robot Do: Human Demo Video to Robot Action Plan via Vision Language Model

### v1

@@ IV-A Task Design
We are particularly interested in the long-horizon daily tasks and construction tasks that can be decomposed into a series of pick-and-place sub-tasks for their clear temporal order and easy demonstration. As illastratued in Figure 3, we collected a set of human demonstration videos covering three diverse categories as the evaluation tasks:
@@
Wooden Block Stacking Task. We also collected demonstrations of a human playing with wooden blocks to simulate a block-building game-play scenario or a miniature construction setup. The key feature of this category is that the visual appearance of the objects is highly similar, which requires relatively precise reasoning of spatial relations, creating a well-known challenging case for the current vision language models. We show that SeeDo can overcome this challenge with the help of the visual prompts from the Visual Perception module. In total, this task contains 39 demonstrations.
@@ IV-B Evaluation Metrics
Specifically, a long-horizon pick-and-place task can be decomposed into a series of single pick-and-place sub-task steps arranged sequentially over time. Each pick-and-place action establishes a relative spatial relation pair between two objects. The final state after the completion of the entire long-horizon task is thus a set of these relational pairs. This provided a measurement of the alignment between the sequence of pick-drop steps generated by a model and those demonstrated in the human videos. Based on this, we propose three metrics: Task Success Rate (TSR), Final-state Success Rate (FSR), and Step Success Rate (SSR), to evaluate the completeness of the generated plan.
@@
  • Spatial Error occurs when the objects are correctly identified and distinguished, but there is an error in reasoning about the spatial relations between them, reflecting the spatial reasoning capability.
@@ IV-C Baselines
Closed-source VLM. We compared SeeDo with the closed-source, commercial model, Gemini 1.5 Pro, which is ranked as the state-of-the-art model for video understanding [49]. We slightly adjusted the prompts since its API can directly take video input. Other than that, the key parts of the prompts are kept consistent with that of SeeDo.
Open-source VLMs. We compared SeeDo with LLaVA One Vision [13] and VILA [45], which also rank top on the video analysis benchmark [49]. We followed their practice and code to uniformly sample 16 frames from each video as the input. Prompts are slightly adjusted to optimize their model’s behavior. The full prompts will be released.
SeeDo variants with GPT-4o. Since we choose GPT-4o as the VLM in SeeDo, we additionally tested three of its variants containing GPT-4o. GPT-4o Init+Final takes only the initial and the final frames of a video as its input. The purpose is to validate whether video input is necessary. To ablate on the Keyframe Selection module, we tested a GPT-4o Unif. which uniformly samples 16 frames without hand detection. As for the Visual Perception module, we compared a GPT-4o without visual prompting. These ablation results are reported respectively in Table II and Table III.
@@ IV-D Results
Main results. Table I presents the overall results of SeeDo and baselines on all three tasks. SeeDo outperforms all closed-source and open-source video VLM baselines across TSR, FSR, and SSR. Since the demonstration videos are all long-horizon, both TSR and FSR require a precise understanding of the entire pick-and-place task from the full-length video. This imposes high accuracy demands on all three modules of SeeDo. Consequently, we observe that when temporal order is incorrectly comprehended, the final state is also wrong. An exception is the GPT-4o Init+Final experiment, where its FSR on the garment task is slightly higher. This shows GPT-4o’s strong reasoning ability. Nevertheless, its overall TSR and SSR are still poor indicating weakness in temporal reasoning. Meanwhile, we observed that even in cases where TSR and FSR fail, SeeDo is often able to successfully interpret most of the task steps in the correct temporal order. As a result, SeeDo’s SSR accuracy exceeds 70 percent for the two daily tasks and 50 percent for the block stacking tasks.
@@
TABLE II: Different sampling. (T/F/S) stands for TSR, FSR, and SSR. I+F stands for initial+final frames. Unif. means uniform samplings. indicates that the context limit is often exceeded.
@@
Abation on the visual prompting for Spatial Reasoning. For pick-and-place task, describing the relative spatial positions between objects is vital. Our wooden block stacking tasks in essence require more precise spatial relationships. To explore the impact of visual prompts on the reasoning capability for complex relative spatial relations, we experimented SeeDo without visual prompt as the baseline to compare with the SeeDo with visual prompting, which uses mask contours, tracking IDs, and the corresponding coordinates of contour centers as prompts. . Experiment results on the wooden block task are shown in Table III. The contrast in the percentage of vision and spatial errors suggests that visual prompting significantly enhances spatial reasoning capabilities.
TABLE III: Ablation on visual prompting (V.P.) on Block Tasks.
@@
Failure Case Analysis. To better understand the performance, we categorized failure cases based on the three error types discussed in Sec. IV-B and calculated their frequency across all demonstrations. As shown in Figure 5, SeeDo consistently has lower error rates across all categories compared to other models, particularly excelling in a notably lower temporal error rate. However, spatial errors remain the main source of SeeDo ’s failures This could be largely attributed to the limited spatial intelligence of current VLMs and the imperfect tracking in the visual perception module. Occasionally, mismatches occur between the two modules, where tracking associates an object with a text description that the VLM refers to as another. This indicates room for future improvement. It is worth noting that the error types are not mutually exclusive and there is a coupling effect. For instance, VILA has a high temporal error rate due to frequently generating erroneous plans with repetitive steps or omitting too many steps. These failure cases are marked as temporal errors, leading to a lower relative incidence of spatial errors. This should not be interpreted as evidence of strong spatial reasoning in VILA, but rather as a consequence of its frequent temporal errors.
Figure 5: Error type percentages of all the failure cases of all the methods. Note that error types are not exclusive. The barplot of the total success rates on all tasks is also presented.
@@ V Conclusion and Limitations
This paper tackles the challenge of extracting robot task plans directly from human demonstration videos using large vision language models (VLMs). We introduce a pipeline, SeeDo, that significantly improves temporal understanding, spatial relationship reasoning, and object differentiation, particularly in cases where objects have similar appearances, outperforming existing video VLMs. Through comprehensive evaluation, SeeDo demonstrates state-of-the-art performance on long-horizon tasks of a series of pick-and-place actions in diverse categories. However, SeeDo still faces many limitations, and we discuss several below.
@@
Limited spatial intelligence. While the visual perception module significantly improves SeeDo’s ability to differentiate left and right spatial relations, it still makes mistakes in tasks requiring more precise spatial reasoning. This is particularly evident in wooden block stacking, where spatial relations in multiple directions are critical for success. We call for future VLMs with stronger spatial intelligence.
Under-defined spatial positioning. Describing the spatial positions of objects is inherently complex. In this work, SeeDo only describes the relative position as a limited set of high-level relative spatial relation pairs. It relies on calling the action primitive to determine the exact positions, which makes it less competent for tasks requiring precise manipulation. In future work, we will explore extracting more precise spatial positioning from the demonstration videos.
@@ V-B Evaluation Metrics
15:  return $\frac{\texttt{MATCH}}{||\mathcal{G}||}$
For all three metrics, we report the average across all demonstrations as the performance of the models.

### v2

@@ IV-A Task Design
We focus on long-horizon daily and construction tasks that decompose naturally into a series of pick-and-place sub-tasks. These tasks represent some common robotics scenarios that feature a clear temporal sequence and dynamic interactions that cannot be adequately captured with still images or brief descriptions. Their reliance on sequential dependencies makes it nearly impossible to specify or replicate the tasks using static visuals or a few words, underscoring the need for full demonstration videos.
As illustrated in Figure 3, we curate a set of human demonstration videos covering three diverse categories as the evaluation tasks:
@@
Wooden Block Stacking Task. We also collected demonstrations of a human playing with wooden blocks to simulate a block-building game-play scenario or a miniature construction site setup. The key feature of this category is that the visual appearance of the objects is highly similar, which requires relatively precise understanding of spatial relations, creating a well-known challenging case for the current vision language models. We show that SeeDo can overcome this challenge with the help of the visual prompts from the Visual Prompting module. In total, this task contains 39 demonstrations.
Our method operates purely on relative spatial relationships (e.g., left, right, above, below) extracted from the demonstration, without relying on fixed camera viewpoints; thus, the execution replicates the same relative positions regardless of whether the input is from a first-person or third-person perspective.
@@ IV-B Evaluation Metrics
Specifically, a long-horizon pick-and-place task can be decomposed into a series of single pick-and-place sub-task steps arranged sequentially over time. Each pick-and-place action establishes a relative spatial relation pair between two objects. The final state after the completion of the entire long-horizon task is thus a set of these relational pairs. This provided a measurement of the alignment between the sequence of pick-drop steps generated by a model and those demonstrated in the human videos. Based on this, we propose three metrics: Task Success Rate (TSR), Final-state Success Rate (FSR), and Step Success Rate (SSR), to evaluate the completeness of the generated task plan.
@@
  • Spatial Error occurs when the objects are correctly identified and distinguished, but there is an error in understanding the spatial relations between them, reflecting the spatial understanding capability.
@@ IV-C Baselines
Video understanding VLMs. We compare SeeDo with VLMs that are capable of direct video understanding. This includes the proprietary Gemini 1.5 Pro, a leading model for video understanding [48], as well as several open-source VLMs that also exhibit strong video understanding abilities, including LLaVA-OneVision [10], LLaVA-NeXT-Video-7B-DPO and VILA [46]. For consistency, we use the same core prompts as SeeDo across all baselines. Gemini 1.5 Pro natively supports video input, while for the open-source models, we follow their official implementations, uniformly sampling 16 frames per video as input. For the open-source baselines, prompts are slightly adjusted to optimize each model’s performance, and the full prompt details will be released with our code.
GPT-4o with alternative frame sampling strategies. Since SeeDo utilizes GPT-4o as its VLM, we further test three variants of GPT-4o using different frame sampling strategies while keeping the same prompts:
  • GPT-4o Init+Final: Uses only the first and last frames of each video to assess whether providing video input is necessary or if start- and end-state images alone suffice
  • GPT-4o Unif.: Following the method in  [63], uniformly samples 16 frames without hand detection, to ablate the Keyframe Selection module.
  • GPT-4o without Visual Prompting: Evaluates the impact of removing the Visual Prompting module.
These two ablation baselines are reported respectively in Table II and Table III.
@@ IV-D Results
Main results. Table I presents the overall results of SeeDo and baselines on all three tasks. SeeDo outperforms all closed-source and open-source video VLM baselines across TSR, FSR, and SSR. Since the demonstration videos are all long-horizon, both TSR and FSR require a precise understanding of the entire pick-and-place task from the full-length video. This imposes high accuracy demands on all three modules of SeeDo. Consequently, we observe that when temporal order is incorrectly comprehended, the final state is also wrong. An exception is the GPT-4o Init+Final experiment, where its FSR on the garment task is slightly higher. This shows GPT-4o’s strong common sense. However, its overall TSR and SSR are still poor, indicating a weakness in temporal reasoning. Meanwhile, we observed that even in cases where TSR and FSR fail, SeeDo is often able to successfully interpret most of the task steps in the correct temporal order. As a result, SeeDo’s SSR accuracy exceeds 70 percent for the two daily tasks and 50 percent for the challenging block stacking tasks.
@@
TABLE II: Different sampling. (T/F/S) stands for TSR, FSR, and SSR. I+F stands for initial+final frames. Unif. stands for uniform frame samplings.
@@
Ablation on the visual prompting for Spatial Understanding. For pick-and-place task, describing the relative spatial positions between objects is crucial. Our wooden block stacking tasks in essence require more precise spatial understanding. To explore the impact of visual prompts on the understanding capability for complex relative spatial relationships, we experimented SeeDo without visual prompt (SeeDo w/o V.P.) as the baseline to compare with the SeeDo with visual prompting, which uses mask contours, tracking IDs, and the corresponding coordinates of contour centers as prompts. Experiment results on the wooden block task are shown in Table III. The contrast in the percentage of vision and spatial errors suggests that visual prompting significantly enhances spatial understanding capabilities.
TABLE III: Ablation on visual prompting (V.P.) on Wooden Block Stacking Task.
@@
Failure Case Analysis. To better understand the performance, we categorized failure cases based on the three error types discussed in Sec. IV-B and calculated their frequency across all demonstrations. As shown in Figure 5, SeeDo consistently has lower error rates across all categories compared to other models, particularly excelling in a notably lower temporal error rate. However, spatial errors remain the main source of SeeDo ’s failures. This could be largely attributed to the limited spatial intelligence of current VLMs and the imperfect tracking in the visual prompting module. Occasionally, mismatches occur between the two modules, where tracking associates an object with a text description that the VLM refers to as another. This indicates room for future improvement. It is worth noting that the error types are not mutually exclusive and there is a coupling effect. For instance, VILA has a high temporal error rate due to frequently generating erroneous plans with repetitive steps or omitting too many steps. These failure cases are marked as temporal errors, leading to a lower relative incidence of spatial errors. This should not be interpreted as evidence of a strong spatial understanding in VILA, but rather as a consequence of its frequent temporal errors.
Figure 5: Error type percentages of all the failure cases of all the methods. Note that error types are not exclusive. The barplot of the total success rates on all tasks is also presented. LLaVA-OV represents the LLaVA-OneVision model.
IV-E Real-world Experiment Setup
Real-world experiments have demonstrated that SeeDo can manipulate objects in the physical world using an appropriate LMP. The experiment is conducted on a Universal Robots’ UR10e cobot attached with a Robotiq 2F-85 gripper. We use an Intel RealSense 455 stereo camera to acquire depth information. The camera is mounted on a tripod standing across the robot and hand-eye calibrated in an eye-to-hand setup. The gripper is programmed to move horizontally above the object and lower in the z-axis to grasp the object.
Because the action primitives are hard-coded, we find that sometimes inevitably result in some failure cases due to lack of adaptability. For example, in the wood block stacking example, the dropping distance is often too high and would cause the wood block to bounce upon contact, causing task failure in the real world. Such failure cases can be improved in the future by fine-tuning the robot’s action primitives or substituting it with more adaptive and generalizable robot policies. In its current version, SeeDo’s evaluation only measures the accuracy of the action plan and does not take into account the success rate of actual execution.
@@ V Conclusion and Limitations
This paper tackles the challenge of extracting robot task plans directly from human demonstration videos using vision language models (VLMs). We introduce a pipeline, SeeDo, that significantly improves temporal understanding, spatial relationship reasoning, and object differentiation, particularly in cases where objects have similar appearances, outperforming existing video VLMs. Through comprehensive evaluation, SeeDo demonstrates state-of-the-art performance on long-horizon tasks of a series of pick-and-place actions in diverse categories. However, SeeDo still faces many limitations, and we discuss several below.
@@
Limited spatial intelligence. While the visual perception module significantly improves SeeDo’s ability to differentiate left and right spatial relations, it still makes mistakes in tasks requiring more precise spatial reasoning. We call for future VLMs with stronger spatial intelligence.
Under-defined spatial positioning. Describing the spatial positions of objects is inherently complex. In this work, SeeDo only describes the relative position as a limited set of high-level relative spatial relation pairs, which makes it less competent for tasks requiring precise manipulation. In future work, we will explore extracting more precise spatial positioning from the demonstration videos.
@@ V-B Evaluation Metrics
[…]
For all three metrics, we report the average across all demonstrations as the performance of the models. Notably, we provide an automated evaluation script in our open-source code, which significantly reduces the need for time-consuming manual verification.
Automated Evaluation with LLM. To validate the reliability of the automated evaluation script, we computed pointwise differences between the script-generated evaluation scores and the manually evaluated scores for each data point. We then visualized these differences using kernel density estimation (KDE) plots and reported their mean values in Fig 6. The mean differences for vegetable organization, garment organization, and wooden block organization are 0.084, 0.167, and 0.010, respectively, with an overall mean difference of 0.074 across all tasks. This indicates that, from a statistical perspective, the automated script’s measurements are close to those of manual evaluation. Garment organization exhibits a higher mean difference, likely due to the task’s higher complexity and the prevalence of ambiguous descriptions for visually similar garments, which are prone to mismatches with the ground truth. In vegetable organization, garment organization, and wooden block stacking, the automated script produced results identical to manual evaluation for 27/38, 6/30, and 26/39 subtasks, respectively. These results collectively demonstrate that the automated evaluation script is statistically reliable and suitable for practical use.
Figure 6: KDE plots of pointwise differences between the automated evaluation script and manually verified ground truth across three tasks and the overall. The vertical lines indicate the mean of the differences on different tasks.